OpenStudy (nastech):

determine if the series below converges or diverges. for convergent series, find the sum of the series ∑_(k=0)^∞▒3(1/5) ^k

OpenStudy (anonymous):

hi. are you still there?

OpenStudy (anonymous):

if i am reading this right, it looks like a geometric series

OpenStudy (turingtest):

\[\sum_{k=0}^\infty3(\frac15)^k\]it looks like

OpenStudy (turingtest):

\[\sum_{k=0}^\infty a(n)^k=\frac a{1-n}~~~~~~~\forall n<1\]

OpenStudy (nastech):

I have no idea

OpenStudy (anonymous):

so the first term of your series is 3 because 3(1/5)^0=3

OpenStudy (anonymous):

1/5 is the common ratio

OpenStudy (anonymous):

the sum of the series is given by S=a/(1-r), a is the first term and r is the common ratio

OpenStudy (anonymous):

do this make sense to you?

OpenStudy (turingtest):

oh yeah, my formula is a bit off

OpenStudy (nastech):

i'm trying to put the pieces together

OpenStudy (turingtest):

oh no, it's right, nvm

OpenStudy (anonymous):

questions are ok here.

OpenStudy (nastech):

I get it now, thanks a lot. I'm actually supposed to create a matlab program to solve the question

OpenStudy (anonymous):

ahhh. gotcha. well, now you have the solution to check your work. take care. this geo series converges. take care

OpenStudy (nastech):

you too take care and may God bless you